OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of sqrt 180 feet and a width of sqrt 125 feet. (3 points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (anonymous):

180x2 + 125x2 = 360 + 250 = 590feet